Consider the distribution system given in Problem 7 and assume that most transformers are shielded from lightning to some extent by buildings or trees. Therefore, use a “reasonable estimate” of 70% for the average shielding factor. It is assumed that the total expected number of strikes to the transformer poles is 174 strikes/year and that 5% of lightning return strokes exceed 100 kA. Determine the following:
(a) The total number of strikes per year to the transformers if shielding is taken into consideration
(b) The total number of expected transformer failures per year due to lightning
(c) The annual expected transformer failure rate due to lightning
Answer #1

(a)

answer = (70/100)*174 = 121.8 strikes per year

(b)

answer = (5/100) * 121.8 = 6.09 failures per year

(c)

answer = 6.09/174 = .035

in percentage = .035*100 = 3.5 %